Parker Plastic, Inc., manufactures plastic mats to use with rolling office chairs. Its standard cost information for last year follows:

Standard Quantity Standard Price (Rate) Standard Unit Cost
Direct materials (plastic) 8 sq ft. $ 0.63 per sq. ft. $ 5.04
Direct labor 0.26 hr. $ 10.50 per hr. 2.73
Variable manufacturing overhead (based on direct labor hours) 0.26 hr. $ 1.50 per hr. 0.39
Fixed manufacturing overhead $308,040 ÷ 906,000 units) 0.34


Parker Plastic had the following actual results for the past year:

Number of units produced and sold 1,220,000
Number of square feet of plastic used 11,200,000
Cost of plastic purchased and used $ 6,720,000
Number of labor hours worked 305,000
Direct labor cost $ 3,233,000
Variable overhead cost $ 540,000
Fixed overhead cost $ 361,000


Required:
Calculate Parker Plastic’s direct labor rate and efficiency variances. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable.)
Answer #1

Labor rate variance = (SR-AR) *AH

= (10.50 - 3,233,000/305,000)*305,000

= 30,500 U

Labor efficiency variance = (SH-AH) *SR

= (1,220,000*0.26 - 305,000)*10.50

=128,100 F